an object propelled upwards with an acceleration of 2.0 m / s ^ 2 is
launched from rest. After 6 seconds the fuel runs out. Determine the
speed at this time and the maximum height at which it reaches.
Answer #1

Here we will use

V= u+at

V- final speed , u- initial speed , a- acceleration and t- time

So speed after 6 sec

V=0+2×6= 12 m/s. ANWER

Height reach you n this time

h​​​​​​1​​ = m

Now it will start to deaccelerate with g= 9.8 m/s​​​​​​2​​​​​s2

So height obtain in this period ( using Newton's third equation of motion)

Maximum height = 36+7.35= 43.35 m. Answer